OpenStudy (anonymous):

HELP! A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (da_scienceman):

ok two cases : minimum: -(x-0.525) =0.01? maximum = x-0.525=0.01? U can solve for x in both cases, I guess!

OpenStudy (da_scienceman):

or d other way round? max and min= pls verify!